# Understanding Blockchain Governance: Analyzing Decentralized Voting to Amend DeFi Smart Contracts

### v2

@@ Abstract.
Governance protocols define the means for amending or changing smart contracts without any centralized authority. They distribute the decision-making power to every user of the smart contract: Users vote on accepting or rejecting every change. In this work, we review and characterize decentralized governance in practice, using Compound and Uniswap—two widely used governance protocols—as a case study. We reveal a high concentration of voting power in both Compound and Uniswap: 10 voters hold together 57.86% and 44.72% of the voting power, respectively. Although proposals to change or amend the protocol receive, on average, a substantial number of votes (i.e., 89.39%) in favor within the Compound protocol, they require fewer than three voters to obtain 50% or more votes. We show that voting on Compound proposals can be unfairly expensive for small token holders, and we discover voting coalitions that can further marginalize these users.
@@ 1. Introduction
The transparent decentralized governance of DeFi smart contracts stands in sharp contrast to opaque, centralized governance of crypto exchanges such as FTX (Ian Allison, 2022; Nathan Reiff, 2023). In the case of FTX, opaque decision-making by a small group of individuals resulted in a collapse in trust. This debacle, among others such as the crashes of Silicon Valley Bank (Edward Helmore, 2023) and Luna (Krisztian Sandor and Ekin Genç, 2022; Q.ai, 2022), highlights the need for a mechanism that will allow every stakeholder to participate in critical decisions and engage in discussions in a transparent manner. The use of governance mechanisms by themselves does not guarantee, nevertheless, that all the potentially diverse interests of stakeholders are taken into account in deciding whether to approve any given amendment. Of crucial importance are how governance tokens (i.e., voting power) are distributed among the stakeholders, whether the token holders actively participate in the voting on amendments in practice, how controversial the amendment proposals are, and whether stakeholders form voting coalitions that result in marginalizing some voters. In practice, unequal or unfair distribution of governance tokens undermine the core objective of decentralization, potentially compromising the security of the governance protocol (Haig, 2022; Sam Kessler, 2022).
@@
Next, we summarize our contributions. First, we characterize the Compound and Uniswap protocols’ on-chain voting process, showing that they are active and regularly used, with a steady flow of proposals. The majority of the proposals receive significant support in Compound: on average, $89.39$% of votes are in favor. Then, we reveal a substantial variation in voting costs, from \$$0.03$ to \$$294.02$, with an average of \$7.887.887.887.88111All costs are in US dollars, taking into account the exchange rate at the time of casting the vote. If we normalize the costs per vote by the count of tokens held by users, we obtain an average cost per vote unit of \$$358.54$. Voting costs can, hence, be unfairly expensive for small token holders, which has fairness implications for the decision-making process. We also show that a small group of 10 voters holds a significant amount of voting power ($57.86$% and $44.72$% of all tokens for Compound and Uniswap, respectively) and that proposals only required an average of $2.84$ voters to obtain at least $50$% of the votes. These observations strongly suggest that the voting outcomes may not reflect the preferences of the broader community. Finally, we discover potential voting coalitions among the top voters, which could further exacerbate concerns of voting concentration.
@@ 2.1. Voting Modalities
Off-Chain Voting  This system conducts voting on an off-chain third-party platform and, as a consequence, also establishes the rules for voting, aggregating votes, and determining the results off-chain. Protocols such as Balancer (Balancer.fi, 2023) and Convex Finance (Convex, 2023a), for instance, use Snapshot (Labs, 2023b) for off-chain voting. Snapshot stores the voting data on a P2P network called InterPlanetary File System (IPFS) (Labs, 2023a). The voting process does not require voters to pay any fees and (unlike on-chain voting) promotes participation across all participants, regardless of their level of participation or investment in the protocol. After the voting, this system uses a multi-signature contract to enact the off-chain voting outcome on the blockchain. Typically, an n-of-m multisig contract requires the transaction to be signed by at least n out of the m “admins” to be executed on-chain. The system trusts the multisig “admins”, who are well-known in the community, to implement the voting outcome on the blockchain truthfully. The admins can also, however, refuse the protocol. In Convex Finance, for instance, the admins can choose not to execute a proposal if they deem it harmful, even if it had received the majority of votes and reached a quorum (Convex, 2023b). On-chain voting systems, in contrast, prevent such manipulation of voting outcomes by one or more individuals (after the voting process), since all governance decisions (e.g., voting and execution) happen on the chain.
Token delegations  Some governance protocols (e.g., Compound, Uniswap, and AAVE) require a user to own a certain amount of governance tokens for casting a vote. Users must delegate their tokens either to themselves or, if they do not wish to vote, delegate them to others. The ability to delegate voting power to others facilitates a form of liquid democracy; the token holder who delegates or sells their tokens to another loses their voting power. Delegations allow anyone to buy (or sell) tokens and gain (or lose) voting power instantly. Justin Sun, the founder of (stablecoin) TrueUSD (TrueUSD, 2023), for instance, allegedly borrowed COMP tokens to create and vote for Compound proposal #84, resulting in a governance attack (Thurman, 2022); this proposal was, however, defeated. We discuss attacks on governance in details in §2.3.
@@
We refer the reader to §F (particularly Tab. 3) for a characterization of the voting methods, delegation approaches, and proposal executions in various other governance protocols.
2.2. Real-world decision-making using Compound Governance
Interestingly, Compound has also been utilized for real-world decision-making purposes, such as allocating grants to contributors or hiring an audit company to review the governance protocol through the Compound code (Sukernik, 2021). For instance, on September 29, a bug was introduced in the Comptroller of the Compound Protocol through proposal #62 that allowed users to claim more COMP tokens than they were entitled to, resulting in a loss of \$50 million worth of COMP tokens (Nick Martitsch, 2021). The Compound community sought to hire, through the Compound governance protocol, a smart contract auditor to audit the protocol (Sukernik, 2021). Three companies, ChainSecurity, OpenZeppelin, and Trail of Bits, posted their business plans for discussion and then created proposals via the Compound Governor. Voters were able to vote for their preferred proposal, and the winning proposal was eventually implemented. The losing proposals would have been cancelled by the community’s multi-signature mechanism after the voting period ended, ensuring only one could pass.
@@ 3. Methodology
Smart-contract events  Smart contracts in Ethereum can generate and dispatch events (or “logs”) for signaling various types of activity (e.g., ERC-20 token transfers or state changes) within the contract. Ethereum’s applications can subscribe to these events, or analyze them later since Ethereum persists the events in the blockchain via the “logs” field of the transaction receipt attribute. We used these logs to filter transactions that triggered specific events, e.g., sending, receiving, or swapping tokens. We also filtered and analyzed transactions that triggered events related to governance protocols tokens to track the evolution of each proposal, including when it was created, when users started voting, and when it was executed or canceled. We refer the reader to §D for low-level implementation details on how we filter and gather transactions that trigger events of interest.
@@
Approval $213220$ $645538$
Standard ERC-20 approval event.
DelegateChanged $12095$ $25492$
Emitted when an account changes its delegate.
DelegateVotesChanged $75820$ $118096$
Emitted when a delegate account’s vote balance changes.
Transfer $1886618$ $3715442$
Emitted when users/holders transfer their tokens to another address.
@@
ProposalCanceled $17$ $6$
Emitted when a proposal is canceled.
ProposalCreated $133$ $29$
Emitted when a new proposal is created.
ProposalExecuted $101$ $14$
Emitted when a proposal is executed in the TimeLock.
ProposalQueued $105$ $14$
Emitted when a proposal is added to the queue in the TimeLock.
VoteCast $9500$ $18084$
Emitted when a vote is cast on a proposal: 00 for against, $1$ for in-favor, and $2$ for abstain.
@@
Inferring wallet address ownership  Since an entity can control multiple wallet addresses in the blockchain, identifying the ownership of these wallets helps in grouping together the accounts that are owned by the same entity. This task is complicated as owners are only identifiable if they choose to voluntarily make their identities public. To address this challenge, we combine wallet ownership information from two widely used data sources: Etherscan (Etherscan, 2023b) and Sybil-List (Sybil, 2023b). The former is a blockchain explorer that helps in identifying the top holders of various cryptocurrencies, and the latter, a Uniswap governance tool for discovering delegates addresses (Sybil, 2023a). It uses cryptographic proofs for verifying wallet addresses voluntarily disclosed by the wallet owner. From these two data sources, we gathered the owners of $3191$ public wallet addresses (refer §E). We used these addresses to infer the owners of $17$ ($51.52$%) of the $33$ unique addresses associated with proposal creation, $114$ ($3.42$%) out of $3335$ proposal voters, and $265$ ($0.13$%) out of $210598$ token holders in Compound. By analyzing the top 10 most influential voters for each proposal, determined by the number of delegated tokens they possessed when casting their vote, we were able to infer the ownership of $67$ ($50.37$%) of these $133$ unique addresses. Finally, as an entity can control more than one address, we grouped the addresses we identified belonging to the same entity together to conduct our analysis.
4. Exploring Governance Data: An In-Depth Analysis
@@
Compound protocol has two versions of its governance contract: Alpha and Bravo. Compound Governor Alpha, the first version of the governance contract, was deployed on Mar. 4, 2020 (block #$9601459$) and was active until Mar. 28, 2021 (block #$12126254$) [^2] The Compound Governor Alpha was deployed at the Ethereum smart contract address 0xc0dA01a04C3f3E0be433606045bB7017A7323E38. The improved version, Compound Governor Bravo, was deployed on Mar. 9, 2021 (block #$12006099$) and has been active since Apr. 14, 2021 (block #$12235671$) [^3] The Compound Governor Bravo was deployed at the Ethereum smart contract address 0xc0Da02939E1441F497fd74F78cE7Decb17B66529. Brave introduced several improvements such as smart-contract upgradability (through proxies), a new option for voters to abstain from voting, and the ability for voters to state the reasons behind their voting choices through text comments attached to on-chain votes. The Bravo contract was proposed in proposal #$42$, and it received $1438679.86$ votes from $59$ voters—all but one vote were in favor of its implementation.
@@ 4.1. Control of Governance Tokens
Distribution of token holding  Initially, $42.15$% of the total Compound supply ($10$ million COMP tokens) was allocated to liquidity mining, $23.95$% to shareholders, $22.46$% to the founders and the Compound team, $7.73$% to the community, and $3.71$% to future team members (CoinGecko, 2023a). The public release of COMP tokens started only after proposal #7 was executed on June 15, 2020. This proposal enabled the continued distribution of COMP tokens to the protocol users over time. At the time of our analysis (Nov. 7, 2022), the $10$ million COMP tokens were distributed among $210573$ accounts. The largest holder is Compound Reservoir with $19.24$% ($1924344.52$) of the tokens followed by Binance ($5.97$% or $397289.78$ tokens) and cComp ($5.73$% or $572723.77$ tokens). The Compound Team holds $2.28$% ($228061.62$) and Compound Timelock $1.84$% ($184258.39$) of the tokens. Due to space constraints, we describe the details of Uniswap in §A.
@@
(a) Tokens in circulation
(b) Tokens delegated
@@
Distribution of token delegation  Delegation is a prerequisite for voting (refer §2.1), and Compound and Uniswap allows its participants to delegate their voting rights to others. This ability enables users to delegate their voting power to individuals who share their interests, and allows participants with less voting power to pool their votes together and have a significant voting impact. Users, however, can only delegate all, not a fraction, of their tokens. The protocol, nevertheless, enforces this limitation at the wallet address level. Users can own multiple wallet addresses and divide their tokens into them, thereby allowing them to delegate a subset of their tokens to others (Fritsch et al., 2022; Amico, 2023). To determine if delegated tokens are held by a few voters, we group together all inferred addresses (as discussed in §3) that belong to the same entity and then count the total number of delegated tokens held by each group. We observe, per Fig. 3, that delegated tokens are concentrated among few voters, and we show the distribution of delegated tokens across several top token holder accounts in Fig. 1(b) and in Fig. 13(b). Out of $4186$ COMP delegatee accounts (or accounts with voting rights), the top 50 ($1.19$%) hold $99.23$% of all delegated tokens, giving them significant decision-making power when voting on proposals. Similarly, in Uniswap, the top $50$ ($0.34$%) hold $94.73$% of all UNI delegated tokens. On Nov. 7, 2022, Polychain Capital held the most COMP delegated tokens, with $12.15$% ($330986.09$) followed by Bain Capital Ventures with $11.85$% ($322763.87$) and a16z with $9.40$% ($256046.13$). These three addresses together held $33.41$% ($909796.10$) of all the $2723123.73$ delegated tokens in our analysis.

### v3

@@ Abstract.
Smart contracts are contractual agreements between participants of a blockchain, who cannot implicitly trust one another. They are software programs that run on top of a blockchain, and we may need to change them from time to time (e.g., to fix bugs or address new use cases). Governance protocols define the means for amending or changing these smart contracts without any centralized authority. They distribute the decision-making power to every user of the smart contract: Users vote on accepting or rejecting every change.
In this work, we review and characterize decentralized governance in practice, using Compound and Uniswap—two widely used governance protocols—as a case study. We reveal a high concentration of voting power in both Compound and Uniswap: 10 voters hold together 57.86% and 44.72% of the voting power, respectively. Although proposals to change or amend the protocol receive, on average, a substantial number of votes (i.e., 89.39%) in favor within the Compound protocol, they require fewer than three voters to obtain 50% or more votes. We show that voting on Compound proposals can be unfairly expensive for small token holders, and we discover voting coalitions that can further marginalize these users.
@@ 1. Introduction
The transparent decentralized governance of DeFi smart contracts stands in sharp contrast to opaque, centralized governance of crypto exchanges such as FTX (Ian Allison, 2022; Nathan Reiff, 2023). In the case of FTX, opaque decision-making by a small group of individuals resulted in a collapse in trust. This debacle, among others such as the crashes of Silicon Valley Bank (Edward Helmore, 2023) and Luna (Krisztian Sandor and Ekin Genç, 2022; Q.ai, 2022), highlights the need for a mechanism that will allow every stakeholder to participate in critical decisions and engage in discussions in a transparent manner. The use of governance mechanisms by themselves does not guarantee, nevertheless, that all the potentially diverse interests of stakeholders are taken into account in deciding whether to approve any given amendment. Of crucial importance are how governance tokens (i.e., voting power) are distributed among the stakeholders, whether the token holders actively participate in the voting on amendments in practice, how controversial the amendment proposals are, and whether stakeholders form voting coalitions that result in marginalizing some voters. In practice, unequal or unfair distribution of governance tokens undermine the core objective of decentralization, potentially compromising the security of the governance protocol (Haig, 2022; Sam Kessler, 2022; Messias et al., 2023).
@@
1.1. Research questions
The challenge of achieving true decentralization in governance protocols is exacerbated by the concentration of voting power. When a minority of participants controls a significant majority of tokens, they can make decisions that primarily serve their interests, often to the detriment of others. Therefore, ensuring an equitable distribution of tokens becomes key in achieving decentralization within these protocols.
Below, we discuss the research questions addressed in this paper.
[…]
This research question aims to investigate the activity level of the these protocols and its community engagement. Specifically, we seek to understand how frequently amendments are proposed and voted on in the protocol. Examining these aspects allows us to assess the level of participation and the extent to which the community actively contributes to improving the protocol.
[…]
This research question aims to investigate the distribution of tokens among its participants. We aim to understand the size of the set of voters who determine outcomes for amendments and assess the decentralization of their tokens. This is crucial for proposing fairness in these protocols based on their token distribution.
[…]
Voting in on-chain governance protocols involves transaction fees that vary based on network congestion. This research question explores the impact of voting costs on voter participation in these protocols. Understanding these costs provides insights into the fairness of the decision-making process and identifies potential barriers that may discourage certain participants from exercising their voting rights. For instance, if the voting cost is too high and voting power is concentrated among a few powerful voters, users may be reluctant to participate in voting. This can be explained from the perception that a handful of powerful voters can easily pass a proposal, making individual less-powerful voter seem less impactful.
[…]
This research question analyzes the voting patterns of delegates in these protocols and investigates whether voters form coalitions. Coalition formation may lead to the marginalization of certain voters, compromising the core principle of decentralization. Understanding the presence and impact of coalition formation provides valuable insights into decision-making dynamics among voters and helps mitigate the concentration of tokens or voting power within these protocols.
1.2. Contributions
More specifically, we summarize our contributions as follows
[…]
@@ 2. Decentralized governance
2.1. Types of Existing Governance Protocols
There are various smart contract applications that utilize decentralized governance protocols for decision-making, including those for lending, decentralized exchanges (DEXes), and stablecoins, among others. An example of such protocols can be found on the Ethereum blockchain, where a number of these applications are available. We have selected some of the most protocols that use decentralized governance for decision-making. Table 1 presents $11$ protocols, including Maker Executive and Maker Pooling, which are part of the MakerDAO (MakerDAO, 2023) stablecoin protocol responsible for the DAI token. These protocols use decentralized governance mechanisms, and we characterize them based on whether their votes are cast on- or off-chain, the delegation methods they use, how they aggregate the votes, and how the proposal outcome take effect.
In our work we focus on two established protocols providing decentralized governance: Compound (Leshner and Hayes, 2019) and Uniswap (Adams et al., 2021). These are two leading governance protocols.
Protocol Type Voting Who can vote? Delegation Voting Aggregation How proposals are implemented?
AAVE (AAVE, 2023) Lending on-chain addresses with delegated tokens yes on-chain on-chain via an SC call.
Balancer (Balancer.fi, 2023) DEX off-chain stakers with locked tokens yes (off-chain) off-chain via 6-of-11 multisig.
Compound (Leshner and Hayes, 2019) Lending on-chain addresses with delegated tokens yes on-chain on-chain via an SC call.
Convex Finance (Convex, 2023a) Yield Farming off-chain stakers with locked tokens yes (off-chain) off-chain via 3-of-5 multisig.
Curve (Curve, 2023) DEX on-chain stakers with locked tokens yes on-chain on-chain through an SC call.
Maker Executive (MakerDAO, 2023) Stablecoin on-chain holders no on-chain New Governance Contract requires more MKR staked than previous.
Maker Polling (MakerDAO, 2023) Stablecoin on-chain addresses with delegated tokens yes off-chain Engineers at Maker create the governance contract based on the voting outcome.
Uniswap (Adams et al., 2021) DEX on-chain addresses with delegated tokens yes on-chain on-chain via an SC call.
1Inch (1inch, 2024) DEX Aggregator off-chain stakers with locked tokens yes off-chain via 7-of-12 multisig.
Ampleforth (Ampleforth, 2024) Stablecoin on-chain addresses with delegated tokens yes on-chain on-chain via an SC call.
Rarible (Rarible, 2024) NFT on-chain stakers with locked tokens yes on-chain on-chain via an SC call. However, a 3-seat security council has veto power over proposals that they deem malicious.
@@ 2.2. Voting Modalities
Off-Chain Voting  This system conducts voting on an off-chain third-party platform and, as a consequence, also establishes the rules for voting, aggregating votes, and determining the results off-chain. Protocols such as Balancer (Balancer.fi, 2023) and Convex Finance (Convex, 2023a), for instance, use Snapshot (Labs, 2023b) for off-chain voting. Snapshot stores the voting data on a P2P network called InterPlanetary File System (IPFS) (Labs, 2023a). The voting process does not require voters to pay any fees and (unlike on-chain voting) promotes participation across all participants, regardless of their level of participation or investment in the protocol. After the voting, this system uses a multi-signature contract to enact the off-chain voting outcome on the blockchain. Typically, an n-of-m multisig contract requires the transaction to be signed by at least n out of the m “admins” to be executed on-chain. The system trusts the multisig “admins”, who are well-known in the community, to implement the voting outcome on the blockchain truthfully. The admins can also, however, refuse the proposal. In Convex Finance, for instance, the admins can choose not to execute a proposal if they deem it harmful, even if it had received the majority of votes and reached a quorum (Convex, 2023b). On-chain voting systems, in contrast, prevent such manipulation of voting outcomes by one or more individuals (after the voting process), since all governance decisions (e.g., voting and execution) happen on the chain.
Token delegations  Some governance protocols (e.g., Compound, Uniswap, and AAVE) require a user to own a certain amount of governance tokens for casting a vote. These tokens can be obtained through airdrops (i.e., given away by the protocol) (Messias et al., 2023), purchased or traded on exchange. Users must delegate their tokens either to themselves or, if they do not wish to vote, delegate them to others. The ability to delegate voting power to others facilitates a form of liquid democracy; the token holder who delegates or sells their tokens to another loses their voting power. Delegations allow anyone to buy (or sell) tokens and gain (or lose) voting power instantly. Justin Sun, the founder of (stablecoin) TrueUSD (TrueUSD, 2023), for instance, allegedly borrowed COMP tokens to create and vote for Compound proposal #84, resulting in a governance attack (Thurman, 2022); this proposal was, however, defeated. We discuss attacks on governance in details in §2.4.
@@
2.3. Real-world decision-making using Decentralized Governance
Interestingly, there are cases where decentralized governance has been applied to real-world decisions. For example, Compound has been utilized for real-world decision-making purposes, such as allocating grants to contributors or hiring an audit company to review the governance protocol through the Compound code (Sukernik, 2021). For instance, on September 29, a bug was introduced in the Comptroller of the Compound Protocol through proposal #62 that allowed users to claim more COMP tokens than they were entitled to, resulting in a loss of \$50 million worth of COMP tokens (Nick Martitsch, 2021; Loewen, 2021). The Compound community sought to hire, through the Compound governance protocol, a smart contract auditor to audit the protocol (Sukernik, 2021). Three companies, ChainSecurity, OpenZeppelin, and Trail of Bits, posted their business plans for discussion and then created proposals via the Compound Governor. Voters were able to vote for their preferred proposal, and the winning proposal was eventually implemented. The losing proposals would have been cancelled by the community’s multi-signature mechanism after the voting period ended, ensuring only one could pass.
@@ 3. Data collection
Smart-contract events  Smart contracts in Ethereum can generate and dispatch events (or “logs”) for signaling various types of activity (e.g., ERC-20 token transfers or state changes) within the contract. We used these logs to filter transactions that triggered specific events, e.g., sending, receiving, or swapping tokens. We also filtered and analyzed transactions that triggered events related to governance protocols tokens to track the evolution of each proposal, including when it was created, when users started voting, and when it was executed or canceled. We refer the reader to §B for low-level implementation details on how we filter and gather transactions that trigger events of interest.
@@
Approval $213220$ $645538$ Standard ERC-20 approval event.
DelegateChanged $12095$ $25492$ Emitted when an account changes its delegate. This means that the delegatee will receive voting power from the sender. Users can only delegate to one address at a time, and the number of votes added to the delegatee’s vote count is equal to the user’s token balance. The delegation of votes will take effect from the current block until the sender either delegates to a different address or transfers their tokens.
DelegateVotesChanged $75820$ $118096$ Emitted when a delegate account’s vote balance changes.
Transfer $1886618$ $3715442$ Emitted when users/holders transfer their tokens to another address.
@@
ProposalCanceled $17$ $6$ Emitted when a proposal is canceled.
ProposalCreated $133$ $29$ Emitted when a new proposal is created.
ProposalExecuted $101$ $14$ Emitted when a proposal is executed in the TimeLock.
ProposalQueued $105$ $14$ Emitted when a proposal is added to the queue in the TimeLock.
VoteCast $9500$ $18084$ Emitted when a vote is cast on a proposal: 00 for against, $1$ for in-favor, and $2$ for abstain.
@@
Inferring wallet address ownership  Since an entity can control multiple wallet addresses in the blockchain (also known as Sybils (Iqbal and Matulevičius, 2021)), identifying the ownership of these wallets helps in grouping together the accounts that are owned by the same entity. This task is complicated as owners are only identifiable if they choose to voluntarily make their identities public. To address this challenge, we combine wallet ownership information from two widely used data sources: Etherscan (Etherscan, 2023b) and Sybil-List (Sybil, 2023b). The former is a blockchain explorer that helps in identifying the top holders of various cryptocurrencies, and the latter, a Uniswap governance tool for discovering delegates addresses (Sybil, 2023a). It uses cryptographic proofs for verifying wallet addresses voluntarily disclosed by the wallet owner. From these two data sources, we gathered the owners of $3191$ public wallet addresses (refer §C). We used these addresses to infer the owners of $17$ ($51.52$%) of the $33$ unique addresses associated with proposal creation, $114$ ($3.42$%) out of $3335$ proposal voters, and $265$ ($0.13$%) out of $210598$ token holders in Compound. By analyzing the top 10 most influential voters for each proposal, determined by the number of delegated tokens they possessed when casting their vote, we were able to infer the ownership of $67$ ($50.37$%) of these $133$ unique addresses. Finally, as an entity can control more than one address, we grouped the addresses we identified belonging to the same entity together to conduct our analysis.
4. An analysis of Compound’s governance
@@
Compound protocol has two versions of its governance contract: Alpha [^2] The Compound Governor Alpha was deployed at the Ethereum smart contract address 0xc0dA01a04C3f3E0be433606045bB7017A7323E38. and Bravo. [^3] The Compound Governor Bravo was deployed at the Ethereum smart contract address 0xc0Da02939E1441F497fd74F78cE7Decb17B66529. Compound Governor Alpha, the first version of the governance contract, was deployed on Mar. 4, 2020 (block #$9601459$) and was active until Mar. 28, 2021 (block #$12126254$). The improved version, Compound Governor Bravo, was deployed on Mar. 9, 2021 (block #$12006099$) and has been active since Apr. 14, 2021 (block #$12235671$). Brave introduced several improvements such as smart-contract upgradability (through proxies), a new option for voters to abstain from voting, and the ability for voters to state the reasons behind their voting choices through text comments attached to on-chain votes. The Bravo contract was proposed in proposal #$42$, and it received $1438679.86$ votes from $59$ voters—all but one vote were in favor of its implementation (Labs, 2021).
@@ 4.1. Control of Governance Tokens
Distribution of token holding  Initially, $42.15$% of the total Compound supply ($10$ million COMP tokens) was allocated to liquidity mining, $23.95$% to shareholders, $22.46$% to the founders and the Compound team, $7.73$% to the community, and $3.71$% to future team members (CoinGecko, 2023a). The public release of COMP tokens started only after proposal #7 was executed on June 15, 2020 (Labs, 2020). This proposal enabled the continued distribution of COMP tokens to the protocol users over time. At the time of our analysis (Nov. 7, 2022), the $10$ million COMP tokens were distributed among $210573$ accounts. The largest holder is Compound Reservoir with $19.24$% ($1924344.52$) of the tokens followed by Binance ($5.97$% or $397289.78$ tokens) and cComp ($5.73$% or $572723.77$ tokens). The Compound Team holds $2.28$% ($228061.62$) and Compound Timelock $1.84$% ($184258.39$) of the tokens. We discuss further details of Uniswap in §5.
@@
(a) COMP tokens in circulation
(b) COMP tokens delegated
@@
4.1.1. Distribution of token delegation
Delegation is a prerequisite for voting (refer §2.2), and Compound and Uniswap allows its participants to delegate their voting rights to others. This ability enables users to delegate their voting power to individuals who share their interests, and allows participants with less voting power to pool their votes together and have a significant voting impact. Users, however, can only delegate all, not a fraction, of their tokens. The protocol, nevertheless, enforces this limitation at the wallet address level. Users can own multiple wallet addresses and divide their tokens into them, thereby allowing them to delegate a subset of their tokens to others (Fritsch et al., 2022; Amico, 2023). To determine if delegated tokens are held by a few voters, we group together all inferred addresses (as discussed in §3) that belong to the same entity and then count the total number of delegated tokens held by each group. We observe, per Fig. 3, that delegated tokens are concentrated among few voters, and we show the distribution of delegated tokens across several top token holder accounts in Fig. 2(b) and in Fig. 15(b). Out of $4186$ COMP delegatee accounts (or accounts with voting rights), the top 50 ($1.19$%) hold $99.23$% of all delegated tokens, giving them significant decision-making power when voting on proposals. Similarly, in Uniswap, the top $50$ ($0.34$%) hold $94.73$% of all UNI delegated tokens. On Nov. 7, 2022, Polychain Capital held the most COMP delegated tokens, with $12.15$% ($330986.09$) followed by Bain Capital Ventures with $11.85$% ($322763.87$) and a16z with $9.40$% ($256046.13$). These three addresses together held $33.41$% ($909796.10$) of all the $2723123.73$ delegated tokens in our analysis.